# RSNet: A Light Framework for The Detection of Multi-scale Remote Sensing Targets

### v1

@@ Abstract
Recent developments in synthetic aperture radar (SAR) ship detection have seen deep learning techniques achieve remarkable progress in accuracy and speed. However, the detection of small targets against complex backgrounds remains a significant challenge. To tackle these difficulties, this letter presents RSNet, a lightweight framework aimed at enhancing ship detection capabilities in SAR imagery. RSNet features the Waveletpool-ContextGuided (WCG) backbone for enhanced accuracy with fewer parameters, and the Waveletpool-StarFusion (WSF) head for efficient parameter reduction. Additionally, a Lightweight-Shared (LS) module minimizes the detection head’s parameter load. Experiments on the SAR Ship Detection Dataset (SSDD) and High-Resolution SAR Image Dataset (HRSID) demonstrate that RSNet achieves a strong balance between lightweight design and detection performance, surpassing many state-of-the-art detectors, reaching 72.5% and 67.6% in respectively with 1.49M parameters. Our code will be released soon.
@@
As SAR image data scale and complexity increase, traditional image processing methods are becoming insufficient [9]. Recent advances in convolutional neural networks (CNNs) and Transformer architectures have significantly improved target detection [10], [11]. These deep learning methods are generally categorized into candidate region-based and regression-based approaches, with the latter excelling in real-time detection by directly regressing targets at multiple scales. For airborne SAR detection, one-stage methods strike a promising balance between speed and accuracy; however, due to the computational and energy limitations of airborne platforms, a lightweight model framework is essential for practical deployment.
@@
To tackle challenges in remote sensing image detection, we present RSNet, an optimized network based on YOLOv8, featuring improvements to its backbone, neck, and head. The backbone incorporates the Waveletpool-ContextGuided (WCG) structure, while the head employs the Waveletpool-StarFusion (WSF) network. Additionally, a Lightweight-Shared (LS) module in the detection head reduces the model’s parameter load.
@@
We developed the Waveletpool-ContextGuided (WCG) architecture, enhancing detection accuracy through anti-aliasing and contextual integration, achieving superior performance with fewer parameters.
The Waveletpool-StarFusion (WSF) network employs wavelet upsampling pooling and starnet convolution layers, significantly reducing parameters while maintaining computational efficiency.
We introduced the Lightweight-Shared (LS) module, inspired by groupnorm, to further minimize the detection head’s parameter load, thereby enhancing the overall efficiency of RSNet.
@@ II-A Overall Architecture
This paper introduces RSNet, a lightweight network for ship detection in SAR images, structured into backbone, neck, and head components as illustrated in Figure 1. The backbone processes the input image through the first three layers to produce feature maps, which is then enhanced by the WCConv module using multiple scales of wavelet pooling to generate multi-scale feature maps (P3, P4, P5). The ContextGuided layer further enriches feature extraction by integrating contextual information. In the neck, the WSF module employs wavelet upsampling to restore spatial resolution and fuse features with the P4 map. The lightweight starnet module incorporates skip connections for deeper feature extraction, improving small object recognition. Finally, the LS Detector module concatenates features to output bounding boxes, class labels, and confidence scores for detected targets.
@@ II-B Waveletpool-ContextGuided(WCG)
In complex backgrounds, including land structures, clutter, and coastlines, current SAR ship detectors face challenges in accurately locating ships. To achieve better performance with fewer parameters, this paper proposes the Waveletpool-ContextGuided Network (WCG).
@@ II-B1 Wavelet Pooling
In the Wavelet Context-Guided Network (WCG), the WaveletPool module plays a critical role, designed to effectively extract multi-scale features from the input feature map through wavelet transforms. The output of this module is defined by the following equations:
where we define the output feature map from C2f as . and are the dimensions of the downsampled output. is the number of output channels of the WaveletPool layer. The filter set used in the WaveletPool operation consists of four different convolutional kernels defined as:
@@ II-B2 Context Guided
Then a depthwise convolution is employed to capture local features , and another depthwise convolution with a dilation rate is applied to capture the surrounding context :
The local features and surrounding features are concatenated, normalized, and activated, generating the global context feature through the global feature extraction module:
@@ II-C Waveletpool-StarFusion(WSF)
Although the WCG module reduces parameters, this paper proposes the Waveletpool-StarFusion Network (WSF) to better balance parameters and model accuracy.
WaveletUnPool and StarNet enhance SAR ship detection by restoring spatial resolution and retaining rich features for improved contour recognition. StarNet utilizes depthwise separable convolutions and efficient multiplication operations for real-time applicability while effectively mitigating noise. Together, they significantly enhance performance and efficiency in ship detection.
@@ II-C1 WaveletUnPool
WaveletUnPool, in contrast, aims to restore the spatial resolution of the downsampled feature map, effectively reconstructing its original dimensions. It utilizes transpose convolution with a stride of 2, employing the same wavelet filters as described in Equation 2 to upsample the feature map back to a higher resolution.
II-C2 StarNet
StarNet [17] is an advanced deep learning architecture that enhances model representation through Star Operation. By fusing features from different subspaces with element-wise multiplication at each layer, it expands implicit feature dimensions without significantly increasing parameters. Figure 2 illustrates this process in the neck section, showing the initial concatenation of the feature map before inputting it into StarNet, along with the computational flow involved.
@@ II-C2 StarNet
(12) The number of channels in and are both mlp_ratio , with mlp_ratio = 3.
@@ II-D LS Detector
The FOCS [18] shows that groupnorm enhances detection head performance in localization and classification. Shared convolutions reduce model parameters, facilitating lightweight design for resource-constrained devices. To address inconsistencies in handling targets of varying scales, we introduce a scale layer for feature adjustment. The specific structure is shown in Fig 3. These strategies lower parameters and computational load while maintaining accuracy.
@@ III-A Datasets
To evaluate the performance of our model, we utilized two publicly available SAR remote sensing datasets: SSDD [19] and HRSID [20]. The SSDD dataset consists of 1160 images, with 928 allocated for training and 232 for testing, as outlined in the original publication. In the case of the HRSID dataset, we adhered to the official partitioning, employing 3642 images for training and 1962 for testing purposes.
@@ III-C Evaluation Metrics
To evaluate the proposed model’s performance, we used mean average precision at IoU = 0.5 () and IoU = 0.5:0.05:0.95 (), along with the number of parameters (Params) and giga floating-point operations per second (GFLOPs). The mAP metrics assess detection accuracy, while Params and GFLOPs indicate parameters and computational complexity.
@@ III-D Ablation Experiments
To validate the effectiveness of the module, we conducted ablation experiments on the HRSID dataset, as shown in Table II. Then, to further validate the model’s ability to perceive real targets, we conducted feature visualization experiments, and the results are shown in Fig. 4, where our model focuses on observing real targets while minimizes the influence of the surrounding noise.
@@ III-E Comparison Experiments
We selected several representative object detectors, including the classic two-stage Faster R-CNN-FPN, one-stage RetinaNet, and YOLOv6n, YOLOv7-tiny, YOLOv8n and YOLOv11n. Additionally, we compared algorithms specifically designed for target detection in SAR images. The experimental results, presented in Tables III and IV, demonstrate that our algorithm effectively balances parameter quantity and performance, outperforming other methods.
@@ IV CONCLUSIONs
This paper introduces RSNet, a lightweight network for SAR ship detection that enhances accuracy and efficiency through innovative architecture. The WCG structure improves performance with fewer parameters, while the WSF network and LS module further optimize parameters. Future work will explore RSNet’s application to other detection tasks and datasets, expanding its utility in remote sensing.

### v4

@@ Abstract
Recent advancements in synthetic aperture radar (SAR) ship detection using deep learning have significantly improved accuracy and speed, yet effectively detecting small objects in complex backgrounds with fewer parameters remains a challenge. This letter introduces RSNet, a lightweight framework constructed to enhance ship detection in SAR imagery. To ensure accuracy with fewer parameters, we proposed Waveletpool-ContextGuided (WCG) as its backbone, guiding global context understanding through multi-scale wavelet features for effective detection in complex scenes. Additionally, Waveletpool-StarFusion (WSF) is introduced as the neck, employing a residual wavelet element-wise multiplication structure to achieve higher dimensional nonlinear features without increasing network width. The Lightweight-Shared (LS) module is designed as detect components to achieve efficient detection through lightweight shared convolutional structure and multi-format compatibility. Experiments on the SAR Ship Detection Dataset (SSDD) and High-Resolution SAR Image Dataset (HRSID) demonstrate that RSNet achieves a strong balance between lightweight design and detection performance, surpassing many state-of-the-art detectors, reaching 72.5% and 67.6% in respectively with 1.49M parameters. Our code will be released soon.
@@
With SAR image data scale and complexity increasing, traditional image processing methods are becoming insufficient [9]. Recent advances in convolutional neural networks (CNNs) and Transformer architectures have significantly improved target detection [10], [11]. These deep learning methods are generally categorized into candidate region-based and regression-based approaches, with the latter excelling in real-time detection by directly regressing targets at multiple scales. For SAR object detection, one-stage methods strike a promising balance between speed and accuracy; however, due to the computational and energy limitations of airborne platforms, a lightweight model framework is essential for practical deployment.
@@
To tackle challenges in remote sensing image detection, we present RSNet (Restar the Sky Network), an optimized network based on YOLOv8, featuring improvements to its backbone, neck, and head. The backbone incorporates the Waveletpool-ContextGuided (WCG) structure, while the head employs the Waveletpool-StarFusion (WSF) network. Additionally, a Lightweight-Shared (LS) module in the detection head reduces the model’s parameter load.
@@
We developed the WCG architecture, enhancing detection accuracy through anti-aliasing and contextual integration, achieving superior performance with fewer parameters.
The WSF network is constructed by combining wavelet upsampling pooling and element-wise multiplication through the residual structure, significantly reducing parameters while maintaining computational efficiency.
We introduced the LS module using the group normalization strategy to further minimize the parameter burden of the detection head, thereby enhancing the overall efficiency of RSNet.
@@ II-A Overall Architecture
This letter introduces RSNet (Restar the Sky Network), a lightweight network for ship detection in SAR images, structured into backbone, neck, and detect components as illustrated in Fig. 1. The backbone processes the input image through the first three layers to produce feature maps, which is then enhanced by multiple scales of wavelet pooling to generate multi-scale feature maps. The ContextGuided block further enriches feature extraction by integrating contextual information. In the neck, the WSF module employs wavelet upsampling to restore spatial resolution and fuse features. The lightweight star module combines an element-wise multiplication structure with skip connections for deeper feature extraction, enhancing small object recognition capability. Finally, the LS Detector module connects the features to output the bounding boxes, class labels, and confidence scores of the detected objects.
@@ II-B Waveletpool-ContextGuided (WCG)
In complex backgrounds, including land structures, clutter, and coastlines, current SAR ship detectors face challenges in accurately locating ships. To achieve better performance with fewer parameters, this letter proposes WCG.
@@ II-B1 Wavelet Pooling
In the WCG, the WaveletPool module plays a critical role, designed to effectively extract multi-scale features from the input feature map through wavelet transforms. The output of this module is defined by the following formula:
where we define the output feature map from C2f as . is the batch size. and are the dimensions of the downsampled output. is the number of output channels of the WaveletPool block. The filter set used in the WaveletPool operation consists of four different convolutional kernels defined as:
@@ II-B2 Context Guided
where represents the number of the input channels. Then a depthwise convolution is employed to capture local features , and another depthwise convolution with a dilation rate is applied to capture the surrounding features :
The and are concatenated, normalized, and activated, generating the global context feature through the global feature extraction module:
@@ II-C Waveletpool-StarFusion (WSF)
Although the WCG module reduces parameters, this letter proposes WSF to better balance parameters and model accuracy as described in Fig. 2. WaveletUnPool and Star module enhance SAR ship detection by restoring spatial resolution and retaining rich features for improved contour recognition. Star module utilizes depthwise separable convolutions and element-wise multiplication operations for real-time applicability while effectively mitigating noise. Together, they significantly enhance performance and efficiency in ship detection.
@@ II-C1 WaveletUnPool
WaveletUnPool, in contrast, aims to restore the spatial resolution of the downsampled feature map, effectively reconstructing its original dimensions. It utilizes transpose convolution with a stride of 2, employing the same wavelet filters as described in formula 2 to upsample the feature map back to a higher resolution.
II-C2 Star
Element-wise multiplication structure inside StarNet [17] which is an advanced deep learning architecture that enhances model representation. By fusing features from different subspaces with element-wise multiplication at each layer, it expands implicit feature dimensions without significantly increasing parameters. Fig. 2 illustrates this process in the neck section, showing the initial concatenation of the feature map before inputting it into Star module, along with the computational flow involved.
@@ II-C2 Star
(12) -
@@ II-D LS Detector
The FOCS [18] shows that groupnorm enhances detection head performance in localization and classification. Shared convolutions reduce model parameters, facilitating lightweight design for resource-constrained devices. To address inconsistencies in handling targets of varying scales, we introduce a scale layer for feature adjustment. The specific structure is shown in Fig. 3. These strategies lower parameters and computational load while maintaining accuracy.
@@ III-A Datasets
To evaluate the performance of our model, we utilized two publicly available SAR remote sensing datasets: SSDD [19] and HRSID [20]. The SSDD consists of 1160 images, with 928 allocated for training and 232 for testing, as outlined in the original publication. In the case of the HRSID, we adhered to the official partitioning, employing 3642 images for training and 1962 for testing purposes.
@@ III-C Evaluation Metrics
To evaluate the proposed model’s performance, we used mean average precision at IoU = 0.5 () and IoU = 0.5:0.05:0.95 (), along with the number of parameters (Params) and floating-point operations per second (FLOPs). The mAP metrics assess detection accuracy, while Params and FLOPs indicate parameters and computational complexity.
@@ III-D Ablation Experiments
To validate the module’s effectiveness, we conducted ablation experiments on the HRSID dataset, as shown in Table II. We also performed feature visualization experiments, illustrated in Fig. 4, which demonstrate that our model effectively identifies real targets while reducing the influence of surrounding noise.
@@ III-E Comparison Experiments
We selected several representative object detectors, including the classic two-stage Faster R-CNN-FPN [27], one-stage RetinaNet [18], and YOLOv6n [28], YOLOv7-tiny [29], YOLOv8n and YOLOv11n. Additionally, we compared algorithms specifically designed for target detection in SAR images. The experimental results, presented in Tables III and IV, demonstrate that our method effectively balances parameter quantity and performance, outperforming other methods.
@@ IV CONCLUSIONs
This letter introduces RSNet, a lightweight network for SAR ship detection that enhances accuracy and efficiency through its innovative architecture. The WCG backbone improves performance with fewer parameters, facilitating effective feature extraction and context understanding. The WSF neck optimizes feature integration, while the LS module ensures efficient detection with minimal computational overhead. Future work will explore RSNet’s application to other detection tasks and datasets, further expanding its utility in remote sensing.